# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ Abstract
We consider a system of NLS with cubic interactions arising in nonlinear optics without Galilean symmetry. The absence of Galilean symmetry can lead to many difficulties, such as global existence and blowup problems; see [Comm. Partial Differential Equations 46, 11 (2021), 2134–2170]. In this paper, we mainly focus on the influence of the absence of this symmetry on the traveling waves of the NLS system. Firstly, we obtain the existence of traveling solitary wave solutions that are non-radial and complex-valued. Secondly, using the asymptotic analysis method, when the frequency $\omega$ is sufficiently large, we establish the high frequency limit of the traveling solitary wave solution. Finally, for the mass critical case, we provide a novel condition for the existence of global solutions which is significantly different from the classical. In particular, this new condition breaks the traditional optimal assumption about initial data.
Keywords: Schrödinger system; Galilean symmetry; Traveling solitary waves; High frequency limit; Global existence
@@ 1 Introduction and Main Results
In this paper, we consider the Cauchy problem for the following system of nonlinear Schrödinger equations with cubic interaction
$\begin{cases}i\partial_{t}u+\Delta u+\left(\frac{1}{9}|u|^{2}+2|v|^{2}\right)u+\frac{1}{3}\bar{u}^{2}v=0,\\ i\gamma\partial_{t}v+\Delta v+(9|v|^{2}+2|u|^{2})v+\frac{1}{9}u^{3}=0,\end{cases}$ (1.1)
with the initial datum $(u,v)|_{t=0}=(u_{0},v_{0})$. Here $u,v:\mathbb{R}\times\mathbb{R}^{N}\to\mathbb{C}$ and the parameter $\gamma>0$, $\bar{u}$ is the complex conjugate of $u$.
System (1.1) arises from a system of nonlinear Schrödinger equations in a suitable dimensionless form, see [21], where the resonant interaction between the beam of the fundamental frequency and its third harmonic in a diffractive dielectric waveguide is characterized by one transverse dimension. For more details about the physical meaning of this system, one can refer to [5, 22] and the references therein. Models such as (1.1) so-called cascading nonlinear processes. These processes can indeed generate effective higher-order nonlinearities, and they stimulated the study of spatial solitary waves in optical materials with $\chi^{(2)}$ or $\chi^{(3)}$ susceptibilities. Recently, the study of system (1.1) attracted a quantity of attention; the topics cover existence, scattering, blowup, stability, etc. see [19, 18, 8, 1, 6].
Let us review some basic properties of system (1.1). It is well known that the Cauchy problem (1.1) is an infinite-dimensional Hamiltonian system and local well-posedness in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$ (see [3]), which have the following three conservation laws:
@@
$K(f,g)=\|\nabla f\|_{L^{2}}^{2}+\|\nabla g\|_{L^{2}}^{2},$
@@
In this paper, we are interested in the traveling solitary wave solution of system (1.1). Now, we consider traveling solitary wave solutions of (1.1) in the form of
$\left(u_{c,w}(t,x),v_{c,w}(t,x)\right)=\left(e^{i\omega t}\phi(x-ct),e^{3i\omega t}\psi(x-ct)\right).$ (1.6)
@@
$\begin{cases}-\Delta\phi+\omega u=\left(\frac{1}{9}\phi^{2}+2\psi^{2}\right)\phi+\frac{1}{3}\bar{\phi}^{2}\psi,\\ -\Delta\psi+3\gamma\omega u=\left(9\psi^{2}+2\phi^{2}\right)\psi+\frac{1}{9}\phi^{3}.\end{cases}$ (1.7)
The existence of the ground state was obtained by Oliveira and Pastor, see [18]. When $\gamma=3$, one can obtain traveling solitary wave solutions from the standing wave solutions through the Galilean transformation. However, when $\gamma\neq 3$, such a construction does not work due to the lack of Galilean symmetry, Therefore, it is meaningful to study the existence of traveling solitary wave solutions of (1.1) when $\gamma\neq 3$.
The question of the existence and properties of the traveling solitary wave solution of the system (1.1) is widely open. However, in other settings, like the general nonlinear Schrödinger system
$\begin{cases}i\partial_{t}u+\Delta u+\left(a_{1}|u|^{2}+b|v|^{2}\right)u=0,\\ i\partial_{t}v+\Delta v+(a_{2}|v|^{2}+b|u|^{2})v=0,\end{cases}$ (1.8)
standing wave solutions are studied by the variational method. There is a very extensive literature on this subject; we refer, for example, to [24, 23, 20, 14, 16, 4] and references therein. Due to the system (1.8) is invariant under the Galilean transformation. This suggests that one can obtain the desired result by considering the real elliptic system.
However, in this paper, we will focus on the complex elliptic system. For $(\omega,c)\in\mathbb{R}\times\mathbb{R}^{N}$ and $c\neq 0$, (1.6) is a solution of (1.1) if and only if
@@
$\begin{cases}-\Delta\phi+\omega\phi+ic\cdot\nabla\phi-\left(\frac{1}{9}|\phi|^{2}+2|\psi|^{2}\right)\phi-\frac{1}{3}\bar{\phi}^{2}\psi=0,\\ -\Delta\psi+3\gamma\omega\psi+i\gamma c\cdot\nabla\psi-\left(9|\psi|^{2}+2|\phi|^{2}\right)\psi-\frac{1}{9}\phi^{3}=0.\end{cases}$ (1.9)
@@
$S_{\omega,c}(u,v)=$ $E(u,v)+\frac{1}{2}\omega M_{3\gamma}(u,v)+\frac{1}{2}c\cdot P(u,v),$ (1.10)
where $E$, $M_{3\gamma}$ and $P$ are given by (1.3), (1.2) and (1.5), respectively. For the sake of argument, we rewrite the energy functional (1.10) as
$S_{\omega,c}(u,v)=$ $\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i}{2}c\cdot x}u\right)\right|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}$
$+\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i\gamma}{2}c\cdot x}v\right)\right|^{2}+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2}-D(u,v).$
Now we define the function space $X_{\omega,c}$ by
$X_{\omega,c}=\left\{(u,v):\left(e^{-\frac{i}{2}c\cdot x}u,e^{-\frac{i\gamma}{2}c\cdot x}v\right)\in\tilde{X}_{\omega,c}\right\},$
@@
$\tilde{X}_{\omega,c}=H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}),$
@@
$\|(u,v)\|_{\tilde{X}_{\omega,c}}=\|\nabla u\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2}+\omega\|u\|_{L^{2}}^{2}+3\gamma\omega\|v\|_{L^{2}},~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\},~{}\gamma>0.$
Notice that $S_{\omega,c}$ is defined on $X_{\omega,c}$ if
$0\displaystyle 1\leq N\leq 3,~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\},~{}\gamma>$ 0 . (1.11)
This condition comes from the Sobolev embedding, which is used to control the nonlinear term. In this paper, we consider the case $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$ and $N\leq 3$.
@@
$\mathcal{A}_{\omega,c}=\{(\phi,\psi)\in X_{\omega,c}~{}:~{}(\phi,\psi)\neq(0,0),~{}S^{\prime}_{\omega,c}(\phi,\psi)=0\}$
@@
$\mathcal{G}_{\omega,c}=\{(\phi,\psi)\in\mathcal{A}_{\omega,c},~{}S_{\omega,c}(\phi,\psi)\leq S_{\omega,c}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\mathcal{A}_{\omega,c}\}.$
In particular, if $c=0$, $\mathcal{G}_{\omega,0}$ is the set of all ground states.
@@
$\begin{cases}-\Delta\tilde{\phi}+\left(\omega-\frac{|c|^{2}}{4}\right)\tilde{\phi}-\left(\frac{1}{9}|\tilde{\phi}|^{2}+2|\tilde{\psi}|^{2}\right)\tilde{\phi}-\frac{1}{3}e^{i\left(\frac{\gamma}{2}-\frac{3}{2}\right)c\cdot x}\bar{\tilde{\phi}}^{2}\tilde{\psi}=0,\\ -\Delta\tilde{\psi}+\left(3\gamma\omega-\frac{|c|^{2}}{4}\right)\tilde{\psi}-\left(9|\tilde{\psi}|^{2}+2|\tilde{\phi}|^{2}\right)\tilde{\psi}-\frac{1}{9}e^{i\left(\frac{3}{2}-\frac{\gamma}{2}\right)c\cdot x}\tilde{\phi}^{3}=0.\end{cases}$ (1.12)
Now we state our first result.
@@
Let $1\leq N\leq 3$, $\gamma>0$, $\gamma\neq 3$ and $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$.
@@
(ii) If $3\gamma>\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$, then the traveling solitary wave solutions obtained in (i) are non-trivial, that is, $u\neq 0$ and $v\neq 0$.
@@
(i) $N=2$ and $\omega\leq 0$.
(ii) $N=1$ or $N=3$, $\omega\leq-\frac{|c|}{2N}$ with $|c|<\min\left\{2(4-N),\frac{6}{\gamma}(4-N)\right\}$.
@@
$-\Delta Q+3\gamma\omega Q=9Q^{3}.$ (1.13)
@@
$\begin{cases}i\partial_{t}u+\frac{1}{2m}\Delta u=\bar{u}v,\\ i\partial_{t}v+\frac{1}{2M}\Delta u=u^{2},\end{cases}(t,x)\in\mathbb{R}\times\mathbb{R}^{N},~{}~{}m,M>0.$ (1.14)
Fukaya, Hayashi and Inui [9] obtained the existence of traveling solitary wave solutions without mass resonance. In addition, the first author in [12] obtained the similar result for three-wave interaction. In these two cases, the system does not have semi-trivial traveling solitary wave solutions, so the argument is simple. For more about the system with quadratic interaction like (1.14), one can see [7, 11, 17, 10] and the reference therein.
@@
1. When $0<3\gamma\leq\left(\frac{9}{\sqrt{2}}\right)^{\frac{4}{N-4}}$ and $\gamma\neq 3$, the traveling solitary wave solutions may be semi-trivial or non-trivial. This case is left as an interesting open problem.
2. In Theorem 1.1, we assume that $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$ and obtain the existence of traveling solitary waves. In Theorem 1.2, we assume that $\omega\leq 0$ or $\omega\leq-\frac{|c|}{2N}$ and obtain the non-existence of traveling solitary waves. However, for $0<\omega<\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$, $N=2$ or $\frac{|c|}{2N}<\omega<\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}\right\}$, $N=1,3$, the existence or non-existence result is left as an interesting open problem.
@@
Let $1\leq N\leq 3$ and $(u_{\omega},v_{\omega})$ be the nontrivial traveling solitary wave solution of (1.9) (see Theorem 1.1). Assume that $|c|$ is bounded and $\omega\to+\infty$, then we have
[…]
Here $(u_{\infty},v_{\infty})$ is the complex ground state solution of the system (1.9) with $c=0$ and $\omega=1$.
@@
1. In Theorem 1.4, we consider the high frequency limit when the velocity $|c|$ is bounded and the frequency $\omega\to+\infty$. In fact, we can also consider the situation when $|c|\to 0$ and $\omega$ is bounded.
2. Unlike the single equation, see [2, 15], there is no momentum term $P(u,v)$, and it is relatively easy to handle. In our case, things become more difficult since we do not know the sign of the term $c\cdot P(u,v)$ and the functions are all complex-valued.
Finally, we provide an interesting result. This result is the global existence of (1.1) for the $L^{2}$-critical case $(N=2)$. When $N=2$, from [18], it is known that if the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$ and
$M_{3\gamma}(u_{0},v_{0})<M_{3\gamma}(Q,P),$ (1.15)
then the solution of (1.1) is globally, where $(Q,P)$ is the positive ground state solution of the following system:
@@
$D(u,v)\leq C_{opt}^{(1)}K(u,v)M_{3\gamma}(u,v),~{}~{}(u,v)\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2}),$ (1.16)
@@
$C_{opt}^{(1)}=\frac{1}{2}\frac{1}{M_{3\gamma}(Q,P)}.$
We know that the condition (1.15) is sharp in general. Indeed, there exist $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})$ satisfying
$M_{3\gamma}(u_{0},v_{0})=M_{3\gamma}(Q,P),$
such that the system (1.1) has the blowup solution, see [18, Theorem 4.2]. It was also proved that there exists a blowup solution in the weight space $L^{2}(\mathbb{R}^{2},|x|^{2}dx)$, see [18, Theorem 4.7 and Theorem 4.8].
On the other hand, from Appendix B, we have the following sharp Gagliardo-Nirenberg inequality, for $(u,v)\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$,
@@
Obviously, from (1.16) and (1.17), we have $C_{opt}^{(1)}>C_{opt}^{(2)}$, this means that
@@
Let $N=2$, $c\in\mathbb{R}^{2}$, $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})\backslash\{0\}$ and

### v2

# Traveling waves of NLS System arising in optical material without Galilean symmetry

@@ Abstract
We consider a system of NLS with cubic interactions arising in nonlinear optics without Galilean symmetry. The absence of Galilean symmetry can lead to many difficulties, such as global existence and blowup problems; see [Comm. Partial Differential Equations 46, 11 (2021), 2134–2170]. In this paper, we mainly focus on the influence of the absence of this symmetry on the traveling waves of the NLS system. Firstly, we obtain the existence and nonexistence of traveling solitary wave solutions that are non-radial and complex-valued. Secondly, using the asymptotic analysis method, we establish the high frequency limit of the traveling solitary wave solutions when the frequency $\omega$ is sufficiently large. Finally, for the mass critical case, we provide an interesting result for the existence of a global solution that is significantly different from the classical. In particular, the new condition of the initial data breaks the traditional optimal assumption.
Keywords: Schrödinger system; Galilean symmetry; Traveling waves; High frequency limit; Global existence
@@ 1 Introduction and Main Results
In this paper, we consider the Cauchy problem for the following nonlinear Schrödinger (NLS) system with cubic interaction
$\begin{cases}i\partial_{t}u+\Delta u-u+\left(\frac{1}{9}|u|^{2}+2|v|^{2}\right)u+\frac{1}{3}\bar{u}^{2}v=0,\\ i\gamma\partial_{t}v+\Delta v-\alpha v+(9|v|^{2}+2|u|^{2})v+\frac{1}{9}u^{3}=0,\end{cases}$ (1.1)
with the initial datum $(u,v)|_{t=0}=(u_{0},v_{0})$. Here $u,v:\mathbb{R}\times\mathbb{R}^{N}\to\mathbb{C}$ and the parameters $\gamma,\alpha>0$, $\bar{u}$ is the complex conjugate of $u$.
System (1.1) arises from a system of nonlinear Schrödinger equations in a suitable dimensionless form, see [20], where the resonant interaction between the beam of the fundamental frequency and its third harmonic in a diffractive dielectric waveguide is characterized by one transverse dimension. This system is also used to model cascading nonlinear processes. These processes can indeed generate effective higher-order nonlinearities, and they stimulated the study of spatial solitary waves in optical materials with $\chi^{(2)}$ or $\chi^{(3)}$ susceptibilities. For more details about the physical meaning of this system, one can refer to [5, 21] and the references therein. Recently, the study of system (1.1) attracted a quantity of attention in mathematics; the topics cover existence, scattering, blowup, stability, etc. see [18, 17, 24, 1, 6].
Let us review some basic properties of system (1.1). It is well known that the Cauchy problem (1.1) is an infinite-dimensional Hamiltonian system and local well-posedness in $H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N})$, $1\leq N\leq 3$ (see [3]), which have the following three conservation laws:
@@
$K(f,g)=\|\nabla f\|_{L^{2}}^{2}+\|\nabla g\|_{L^{2}}^{2},~{}~{}M_{\alpha}(u(t),v(t))=\|u\|_{L^{2}}^{2}+\alpha\|v\|_{L^{2}}^{2},$
@@
In this paper, we are interested in the traveling solitary wave solution of system (1.1) of the form
$\left(u_{c,w}(t,x),v_{c,w}(t,x)\right)=\left(e^{i(\omega-1)t}\phi(x-ct),e^{3i(\omega-1)t}\psi(x-ct)\right).$ (1.6)
@@
$\begin{cases}-\Delta\phi+\omega\phi=\left(\frac{1}{9}\phi^{2}+2\psi^{2}\right)\phi+\frac{1}{3}\bar{\phi}^{2}\psi,\\ -\Delta\psi+(3\gamma\omega-3\gamma+\alpha)\psi=\left(9\psi^{2}+2\phi^{2}\right)\psi+\frac{1}{9}\phi^{3}.\end{cases}$ (1.7)
The existence of the ground state was obtained by Oliveira and Pastor, see [17]. When $\gamma=3$, one can obtain traveling solitary wave solutions from the standing wave solutions through the Galilean transformation. However, when $\gamma\neq 3$, such construction does not work due to the lack of Galilean symmetry. Therefore, it is interesting to study the existence and nonexistence of traveling solitary wave solutions of (1.1) when $\gamma\neq 3$.
We should mention that in the past 20 years, the following nonlinear Schrödinger system has received more mathematical attention:
$\begin{cases}i\partial_{t}u+\Delta u+\left(a_{1}|u|^{2}+b|v|^{2}\right)u=0,\\ i\partial_{t}v+\Delta v+(a_{2}|v|^{2}+b|u|^{2})v=0,\end{cases}$
This system has good variational frame, and there are many results focusing on standing waves, see[23, 22, 19, 13, 15, 4] and references therein. However, very few results focus on system (1.1), and the nonlinearity term therein seems harder to deal with.
For $(\omega,c)\in\mathbb{R}\times\mathbb{R}^{N}$ and $c\neq 0$, (1.6) is a solution of (1.1) if and only if
@@
$\begin{cases}-\Delta\phi+\omega\phi+ic\cdot\nabla\phi-\left(\frac{1}{9}|\phi|^{2}+2|\psi|^{2}\right)\phi-\frac{1}{3}\bar{\phi}^{2}\psi=0,\\ -\Delta\psi+(3\gamma\omega-3\gamma+\alpha)\psi+i\gamma c\cdot\nabla\psi-\left(9|\psi|^{2}+2|\phi|^{2}\right)\psi-\frac{1}{9}\phi^{3}=0.\end{cases}$ (1.8)
@@
$S_{\omega,c,\alpha-3\gamma}(u,v)=$ $E(u,v)+\frac{1}{2}(\omega-1)M_{3\gamma}(u,v)+\frac{1}{2}c\cdot P(u,v)$
$=$ $\frac{1}{2}K(u,v)+\frac{1}{2}\omega M_{3\gamma}(u,v)+\frac{1}{2}(\alpha-3\gamma)\int|v|^{2}+\frac{1}{2}c\cdot P(u,v)-D(u,v),$ (1.9)
where $E$, $K$, $M_{3\gamma}$ and $P$ are given by (1.3), (1.4), (1.2) and (1.5), respectively. For the sake of argument, we rewrite the energy functional (1) as
$S_{\omega,c,\alpha-3\gamma}(u,v)=$ $\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i}{2}c\cdot x}u\right)\right|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}$
$+\frac{1}{2}\int\left|\nabla\left(e^{-\frac{i\gamma}{2}c\cdot x}v\right)\right|^{2}+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2}+\frac{\alpha-3\gamma}{2}\|v\|_{L^{2}}^{2}-D(u,v).$
Now we define the function space $X_{\omega,c,\alpha-3\gamma}$ by
$X_{\omega,c,\alpha-3\gamma}=\left\{(u,v):\left(e^{-\frac{i}{2}c\cdot x}u,e^{-\frac{i\gamma}{2}c\cdot x}v\right)\in\tilde{X}_{\omega,c,\alpha-3\gamma}\right\},$
@@
$\tilde{X}_{\omega,c,\alpha-3\gamma}=H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}),~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}-\frac{\alpha-3\gamma}{3\gamma}\right\},~{}\gamma>0,$
@@
$\|(u,v)\|_{\tilde{X}_{\omega,c,\alpha-3\gamma}}=\|\nabla u\|_{L^{2}}^{2}+\|\nabla v\|_{L^{2}}^{2}+\omega(\|u\|_{L^{2}}^{2}+3\gamma\|v\|_{L^{2}}^{2})+(\alpha-3\gamma)\|v\|_{L^{2}}^{2}.$
Notice that $S_{\omega,c,\alpha-3\gamma}$ is defined on $X_{\omega,c,\alpha-3\gamma}$ if
$0\displaystyle 1\leq N\leq 3,~{}\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}-\frac{\alpha-3\gamma}{3\gamma}\right\},~{}\gamma>$ 0 . (1.10)
This condition comes from the Sobolev embedding, which is used to control the nonlinear term. In this paper, we consider the case $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}-\frac{\alpha-3\gamma}{3\gamma}\right\}$ and $1\leq N\leq 3$.
@@
$\mathcal{A}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in X_{\omega,c,\alpha-3\gamma}~{}:~{}(\phi,\psi)\neq(0,0),~{}S^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0\}$
@@
$\mathcal{G}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\mathcal{A}_{\omega,c,\alpha-3\gamma},~{}S_{\omega,c,\alpha-3\gamma}(\phi,\psi)\leq S_{\omega,c,\alpha-3\gamma}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\mathcal{A}_{\omega,c,\alpha-3\gamma}\}.$
In particular, if $c=0$, $\mathcal{G}_{\omega,0,\alpha-3\gamma}$ is the set of all ground states.
@@
$\begin{cases}-\Delta\tilde{\phi}+\left(\omega-\frac{|c|^{2}}{4}\right)\tilde{\phi}-\left(\frac{1}{9}|\tilde{\phi}|^{2}+2|\tilde{\psi}|^{2}\right)\tilde{\phi}-\frac{1}{3}e^{i\left(\frac{\gamma}{2}-\frac{3}{2}\right)c\cdot x}\bar{\tilde{\phi}}^{2}\tilde{\psi}=0,\\ -\Delta\tilde{\psi}+\left(3\gamma\omega-3\gamma+\alpha-\frac{\gamma^{2}|c|^{2}}{4}\right)\tilde{\psi}-\left(9|\tilde{\psi}|^{2}+2|\tilde{\phi}|^{2}\right)\tilde{\psi}-\frac{1}{9}e^{i\left(\frac{3}{2}-\frac{\gamma}{2}\right)c\cdot x}\tilde{\phi}^{3}=0.\end{cases}$ (1.11)
Now we state our existence and non-existence results.
@@
Let $1\leq N\leq 3$, $\gamma>0$, $\gamma\neq 3$, and $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}-\frac{\alpha-3\gamma}{3\gamma}\right\}$.
@@
(ii) If $3\gamma\omega+\alpha-3\gamma>\omega{9}^{\frac{4}{4-N}}$ (in particular, if $\alpha=3\gamma$, we only need to assume that $3\gamma>9^{\frac{4}{4-N}}$), then the traveling solitary wave solutions obtained in (i) are non-trivial, that is, $u\neq 0$ and $v\neq 0$.
@@
(i) $N=2$ and $\omega\leq\min\left\{0,\frac{\alpha-3\gamma}{3\gamma}\right\}$.
(ii) $N=1$ or $N=3$, $\omega\leq\min\{-\frac{|c|}{2N},-\frac{|c|}{2N}-\frac{3\gamma-\alpha}{3\gamma}\}$ with $|c|<\min\left\{2(4-N),\frac{6}{\gamma}(4-N)\right\}$.
@@
$-\Delta Q+(3\gamma\omega+\alpha-3\gamma)Q=9Q^{3}.$ (1.12)
@@
$\begin{cases}i\partial_{t}u+\frac{1}{2m}\Delta u=\bar{u}v,\\ i\partial_{t}v+\frac{1}{2M}\Delta v=u^{2},\end{cases}(t,x)\in\mathbb{R}\times\mathbb{R}^{N},~{}~{}m,M>0.$ (1.13)
Fukaya, Hayashi and Inui [8] obtained the existence of traveling solitary wave solutions without mass resonance. In addition, the first author in [11] obtained a similar result for three-wave interactions. In these two cases, the system does not have semi-trivial traveling solitary wave solutions, so the argument is simple. For more about the system with quadratic interaction like (1.13), one can see [7, 10, 16, 9] and the reference therein.
@@
1. When $0<3\gamma\omega+\alpha-3\gamma<\omega{9}^{\frac{4}{4-N}}$ and $\gamma\neq 3$, the traveling solitary wave solutions may be semi-trivial or non-trivial. This case is left as an interesting open problem.
2. In Theorem 1.1, we assume that $\omega>\max\left\{\frac{|c|^{2}}{4},\frac{\gamma|c|^{2}}{12}-\frac{\alpha-3\gamma}{3\gamma}\right\}$ and obtain the existence of traveling solitary waves. In Theorem 1.2, we assume that $\omega\leq 0$ or $\omega\leq\min\{-\frac{|c|}{2N},-\frac{|c|}{2N}-\frac{3\gamma-\alpha}{3\gamma}\}$ and obtain the non-existence of traveling solitary waves. However, we do not know the existence or non-existence result in the remaining case.
@@
Assume $1\leq N\leq 3$. Let $(u_{\omega},v_{\omega})$ be the nontrivial traveling solitary wave solution of (1.8) (see Theorem 1.1). Assume that $|c|$ is bounded and $\omega\to+\infty$, then we have
$\frac{1}{\omega^{\frac{1}{2}}}\left(u_{\omega}\left(\frac{x}{\omega^{\frac{1}{2}}}\right),v_{\omega}\left(\frac{x}{\omega^{\frac{1}{2}}}\right)\right)\to(u_{\infty},v_{\infty})~{}~{}\text{in}~{}~{}H^{1}(\mathbb{R}^{N}).$
Here $(u_{\infty},v_{\infty})$ is the complex ground state solution of the system (1.8) with $c=0$, $\omega=1$ and $\alpha=3\gamma$, that is
$\begin{cases}-\Delta\phi+\phi-\left(\frac{1}{9}|\phi|^{2}+2|\psi|^{2}\right)\phi-\frac{1}{3}\bar{\phi}^{2}\psi=0,\\ -\Delta\psi+3\gamma\psi-\left(9|\psi|^{2}+2|\phi|^{2}\right)\psi-\frac{1}{9}\phi^{3}=0.\end{cases}$ (1.14)
@@
1. In Theorem 1.4, we consider the high frequency limit when the velocity $|c|$ is bounded and the frequency $\omega\to+\infty$. In fact, we can also consider the case when $|c|\to 0$ and $\omega$ is bounded.
2. Unlike the single equation, see [2, 14], there is no momentum term $P(u,v)$. However, in our case, things become more difficult since we do not know the sign of the term $c\cdot P(u,v)$ and the functions are all complex-valued.
Finally, we provide an interesting result. This result is the global existence of (1.1) for the mass-critical case $(N=2)$. When $N=2$, from [17], it is known that if the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$ and
$M_{3\gamma}(u_{0},v_{0})<M_{3\gamma}(Q,P),~{}~{}\alpha=3\gamma,$ (1.15)
then the solution of (1.1) is globally. Here $(Q,P)$ is the positive ground state solution of the following system:
@@
$D(u,v)\leq C_{opt}^{(1)}K(u,v)^{\frac{N}{2}}M_{3\gamma}(u,v)^{2-N/2},~{}~{}(u,v)\in H^{1}(\mathbb{R}^{N})\times H^{1}(\mathbb{R}^{N}),$ (1.16)
@@
$C_{opt}^{(1)}=\frac{(4-N)^{\frac{N-2}{2}}}{N^{\frac{N}{2}}}\frac{1}{M_{3\gamma}(Q,P)}.$
We know that the condition (1.15) is sharp in the case $\gamma=3$. Indeed, there exist $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})$ satisfying
$M_{3\gamma}(u_{0},v_{0})=M_{3\gamma}(Q,P),~{}~{}\alpha=3\gamma=9,$
such that the system (1.1) has the blowup solution, see [17, Theorem 4.2]. It was also proved that there exists a blowup solution in the weight space $L^{2}(\mathbb{R}^{2},|x|^{2}dx)$; see [17, Theorem 4.7 and Theorem 4.8].
Therefore, one question is whether the condition (1.15) is still optimal when $\gamma\neq 3$?
In what follows, we will provide a conclusion (see Theorem 1.5 below) that the condition (1.15) is not sharp in case $\gamma\neq 3$ and $\alpha=3\gamma$.
To show this, we first give the following sharp Gagliardo-Nirenberg inequality (see Appendix B), for $(u,v)\in H^{1}(\mathbb{R}^{2})\times H^{1}(\mathbb{R}^{2})$,
@@
Obviously, from (1.16) and (1.17), we have $C_{opt}^{(1)}>C_{opt}^{(2)}$, which means that
@@
Assume $N=2$, $c\in\mathbb{R}^{2}$ and $\alpha=3\gamma$. Let $u_{0},v_{0}\in H^{1}(\mathbb{R}^{2})\backslash\{0\}$ and